# Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Spatial Reasoning Questions

### v1

@@ 6 Conclusion
Spatial-RAG enhances LLMs’ spatial reasoning by integrating structured spatial retrieval with natural language understanding, bridging the gap between spatial databases and AI-driven question answering. Our framework improves spatial intelligence, enabling applications in urban planning, tourism, navigation, and geographic QA. Extensive evaluations show that Spatial-RAG outperforms existing methods, highlighting its potential to advance spatial analysis, autonomous navigation, and augmented reality. Beyond improving spatial reasoning, Spatial-RAG lays the foundation for future multimodal models that seamlessly integrate textual and spatial data.

### v2

@@ 6 Conclusion
Spatial-RAG enhances LLMs’ spatial reasoning by integrating structured spatial retrieval with natural language understanding, bridging the gap between spatial databases and AI-driven question answering. Our framework improves spatial intelligence, enabling applications in urban planning, tourism, navigation, and geographic QA. Extensive evaluations show that Spatial-RAG outperforms existing methods, highlighting its potential to advance spatial analysis, autonomous navigation, and augmented reality. Beyond improving spatial reasoning, Spatial-RAG lays the foundation for future multimodal models that